OpenStudy (anonymous):

find the solution to the quadratic equation by completing the square. x^2+6x=7

OpenStudy (anonymous):

half of 6 is 3, and 3 squared is 9 go to \[x^2+6x+9=7+9\] or \[(x+3)^2=16\]